# On edge-colouring-games by Erdős, and Bensmail and Mc Inerney

### v1

@@ Abstract
We consider two games proposed by Erdős, and one game by Bensmail and Mc Inerney, all with the same setup of two players alternately colouring one edge of a clique. We give observations and particular behaviour for each of these problems, and prove a first reduction towards confirming the conjecture by Bensmail and Mc Inerney. We state a conjecture for Erdős’ game on the largest induced maximum degree, and extensions to edge-transitive and, respectively, regular graphs.
@@ 1 Introduction
We consider three games: two proposed by Erdős [1] (popularised last year at BCC by Cameron [4] and as problem in the database of Erdős’ problems, https://www.erdosproblems.com/778), and one by Bensmail and Mc Inerney [2], all with a similar setup.
Consider a graph $G$ of order $n$ and size $m.$ Alice and Bob play a game on the edges of $G$, taking turns to colour them red (Alice) and blue (Bob). Let $A$ be the set of red edges and $B$ the set of blue edges at the end. When all edges are coloured, a score is assigned. Here Alice will have $a$ points and Bob have $b$ points, where $a$ and $b$ depend on the final colouring of $G$, i.e., on $G[A]$ and $G[B]$ (the graphs induced by the edges in their colour respectively). Here, $a$ and $b$ will be bounded by the order of the graph, $n$. Both players play optimally, ensuring that their score is as high as possible relative to the other’s score. The score of the game, $s(G)$, equals the difference $a-b$. Alice tries to maximise $s(G)$ and Bob tries to minimise $s(G)$. To ensure a unique outcome $(a,b)$, we assume Alice tries to maximise $a$ for the optimal value of $s=a-b$, and Bob tries to minimise it for the optimal value of $s=a-b$. Equivalently, considering the outcome-function $f(G)=(C+1)a-Cb=C(a-b)+a$ for some $C\geq n$, Alice tries to make $f(G)$ as large as possible and Bob wants $f(G)$ to be as small as possible.
@@
The three games are called the Clique game, the Star game and the vertex-capturing game.
  • In the Clique game, $\operatorname{Clique}$, $a=\omega(G[A])$ and $b=\omega(G[B])$.
  • In the Star game, $\operatorname{Star}$, $a=\Delta(G[A])$ and $b=\Delta(G[B])$.
  • In the vertex-capturing game, $\operatorname{VC}$, at the end of the game; $a$ is the number of vertices for which most incident edges are red, and $b$ is the number of vertices for which most incident edges are blue ($a+b<n$ is possible if there are vertices with half of their incident edges in either colour).
Each game played on a graph $G$ is denoted by $\operatorname{Clique}(G),\operatorname{Star}(G)$ or $\operatorname{VC}(G)$ respectively, when $G=K_{n}$, we write $\operatorname{Clique}(n),\operatorname{Star}(n)$ or $\operatorname{VC}(n).$
The outcome $(a,b)$ and score $a-b$ of the games will be denoted by respectively $\operatorname{out}_{\omega},\operatorname{out}_{\Delta},\operatorname{out}_{\operatorname{VC}}$ and $s_{\omega},s_{\Delta},s_{\operatorname{VC}}$.
By a strategy-stealing argument, one can prove that the first player has an advantage (or at least will never perform worse than the second player). For details, one can read the proof of [2, Thm. 2.1 & 2.2], which applies to all three games.
@@
Erdős conjectured that for $\operatorname{Clique}(n)$, the second player can always force a win as long as $n\not=2,$ and asked the same question for the Star game.
Conjecture 2 ([1]).
$\operatorname{Clique}(n)$ is a player $2$ win, $s_{\omega}(n)=0$, for all $n\in\mathbb{N}\setminus\{2\}.$
@@
Bensmail and Mc Inerney [2, Conj. 7.2] conjectured the exact outcome for the vertex-capturing game (on cliques), and confirmed it for $n\leq 6$, in [2, Prop. 7.6].
@@
$s_{\operatorname{VC}}(n)=\begin{cases}2\mbox{ if }n=2\\ 1\mbox{ if }n\equiv 3\pmod{4}\\ 0\mbox{ otherwise. }\\ \end{cases}$
Despite the interest of many mathematicians, little progress has been made. In our work, we highlight the underlying difficulties that may explain why it remains unresolved.
In Section 2, we consider the three games for Colex graphs. A Colex graph $\mathcal{C}(m)$, where $m=\binom{m_{1}}{2}+m_{2}$ with $0\leq m_{2}<m_{1}$, is the graph formed by adding $m_{2}$ edges between $K_{m_{1}}$ and an additional vertex (if $m_{2}>0$). See Fig. 1 for examples where $(m_{1},m_{2})\in\{(4,3),(5,2)\}.$ As a class of graphs that smoothly extend cliques, this serves as a showcase to challenge certain intuitive expectations by exhibiting particular behaviours. The results are determined by computer code (which implements the idea of Zermelo’s theorem [8]), explained in Section 4.
Having solved some small cases, in Section 3, we extend an observation by [6] using strategy stealing, confirming that $\operatorname{Clique}(n)$ being a win for the first player implies that $\operatorname{Clique}(n+3)$ is a win for the second player. We also prove one implication between the validity of 4 for two residue classes modulo $4$. [^1] This proof by strategy stealing is similar to reductions in [6] and was obtained before their publication, as communicated with them.
@@
If 4 is true for every $n\equiv 3\pmod{4},$ then it is true for every $n\equiv 1\pmod{4}$ as well.
Finally, we give our intuition, formulated in conjectures for the Clique and Star game (stated in the framework of edge-transitive and regular graphs). In particular, we are the first who dare to conjecture the answer for 3.

### v2

@@ Abstract
We study two games proposed by Erdős, and one game by Bensmail and Mc Inerney, all sharing a common setup: two players alternately colour edges of a complete graph, or in the biased version, they colour $p$ and $q$ edges respectively on their turns, aiming to maximise a graph parameter determined by their respective induced subgraphs. In the unbiased case, we give a first reduction towards confirming the conjecture of Bensmail and Mc Inerney, propose a conjecture for Erdős’ game on maximum degree, and extend the clique and maximum-degree versions to edge-transitive and regular graphs. In the biased case, the maximum-degree and vertex-capturing games are resolved, and we prove the clique game with $(p,q)=(1,3)$.
@@ 1 Introduction
We consider three games: two proposed by Erdős [1] (popularised last year at BCC by Cameron [9] and as a problem in the Erdős’ problem database, https://www.erdosproblems.com/778), and one by Bensmail and Mc Inerney [7], all with a similar setup. We first describe the shared principles of these games.
Consider a graph $G$ of order $n$ and size $m.$ Alice and Bob play a game, taking turns to colour the edges of $G$ red (Alice) and blue (Bob), where Alice is the starting player (player $1$), and Bob is the second player (player $2$). When all edges are coloured, a score, which depends on the set of red edges $A$ and the set of blue edges $B$, is assigned. Here Alice has score $a$ and Bob $b$, where $a$ and $b$ depend on the final colouring of $G$, i.e., on $G[A]$ and $G[B]$ (the graphs induced by the edges in their colour respectively). In the studied games, $a$ and $b$ will be bounded by the order of the graph, $n$. Both players play optimally, ensuring that their score is as high as possible relative to the other’s score. The score of the game, $s(G)$, equals the difference aba-b. Alice tries to maximise $s(G)$, while Bob tries to minimise $s(G)$. To ensure a unique outcome $(a,b)$, we assume Alice tries to maximise $a$ for the optimal value of s=abs=a-b, and Bob tries to minimise it for the optimal value of s=abs=a-b. Equivalently, considering the outcome-function f(G)=(C+1)aCb=C(ab)+af(G)=(C+1)a-Cb=C(a-b)+a for some $C\geq n$, Alice wants $f(G)$ to be as large as possible and Bob wants $f(G)$ to be as small as possible.
@@
The three games we will discuss are called the Clique game, the Star game and the vertex-capturing game. When played on a graph $G$, the games are denoted by $\operatorname{Clique}(G),\operatorname{Star}(G)$ or $\operatorname{VC}(G)$ respectively, and when $G=K_{n}$, we write $\operatorname{Clique}(n),\operatorname{Star}(n)$ or $\operatorname{VC}(n).$
  • In the Clique game, $a=\omega(G[A])$ and $b=\omega(G[B])$.
  • In the Star game, $a=\Delta(G[A])$ and $b=\Delta(G[B])$.
  • In the vertex-capturing game, at the end of the game; $a$ is the number of vertices for which most incident edges are red, and $b$ is the number of vertices for which most incident edges are blue ($a+b<n$ is possible if there are vertices with half of their incident edges in either colour).
The outcome $(a,b)$ and score aba-b of the games will be denoted by respectively $\operatorname{out}_{\omega},\operatorname{out}_{\Delta},\operatorname{out}_{\operatorname{VC}}$ and $s_{\omega},s_{\Delta},s_{\operatorname{VC}}$.
By a strategy-stealing argument, one can prove that the first player will never perform worse than the second player. For details, see the proof of [7, Thm. 2.1 & 2.2], which applies to all three games.
@@
Biased (asymmetric) variants of a colouring game were defined by [6], who provided motivation for studying such asymmetric versions. A problem in certain applications can be translated into an asymmetric game (the opponent represents the worst case behaviour).
In the biased setting of the earlier defined games, the first player (Alice) colours $p$ edges at her turn and the second player (Bob) colours $q$ edges at his turn. Here we assume $q>p$ and Bob wins if at the end $b>a$ for the respective parameters of the game, or $q=p$ and Bob wins if $a=b.$
To specify the parameters $p$ and $q$, we write $\operatorname{Clique}_{(p,q)}(G)$ or $\operatorname{Clique}_{(p,q)}(n)$ and use analogous notation for the other two games.
Erdős conjectured that for $\operatorname{Clique}(n)$, the second player can always force a win as long as $n\not=2,$ and asked the same question for $\operatorname{Star}(n)$.
Conjecture 2.
[[1]] $\operatorname{Clique}(n)$ is a player $2$ win, $s_{\omega}(n)=0$, for all $n\in\mathbb{N}\setminus\{2\}.$
@@
Bensmail and Mc Inerney [7, Conj. 7.2] conjectured the exact outcome for the vertex-capturing game (on cliques), and confirmed it for $n\leq 6$, in [7, Prop. 7.6].
@@
[…]
We prove one implication between the validity of ˜4 for two residue classes modulo 44.111This proof by strategy stealing is similar to reductions in [11] and was obtained before their publication, as communicated with them.
@@
If ˜4 is true for every $n\equiv 3\pmod{4}$, then it is true for every $n\equiv 1\pmod{4}$ as well.
This is done in Section˜2, where, for completeness, we also extend an observation by [11] using strategy stealing, confirming that $\operatorname{Clique}(n)$ being a win for the first player implies that $\operatorname{Clique}(n+3)$ is a win for the second player for every $n$.
Finally, we give our intuition, formulated in conjectures for the Clique and Star game (stated in the framework of edge-transitive and regular graphs). In particular, we are the first to conjecture the answer for ˜3.
@@
At the end of their conclusion, Malekshahian and Spiro [11] wondered whether studying $\operatorname{Star}_{(p,q)}(n)$ would be interesting. In Section˜3, more precisely ˜11, we notice that both $\operatorname{Star}_{(p,q)}(n)$ and $\operatorname{VC}_{(p,q)}(n)$ are (up to some small cases) player $2$ wins whenever $q>p.$
In [11], the authors noted that $\operatorname{Clique}_{(p,q)}(n)$ is a player $2$ win for an exponential bound of $q$ in $p$. In Section˜3, we improve this drastically to a (unconditional) quadratic bound and a (conditional) linear bound. Erdős [1] also asked if $\operatorname{Clique}_{(1,2)}{(n)}$ is won by the second player whenever $n\geq 4.$ We conjecture this is indeed the case. Our main result in this direction proves this when $q=2$ is replaced by $q=3$. (Notice that in [11], it is proved with $q=15$)
Theorem 8.
For every $n\geq 4$, $\operatorname{Clique}_{(1,3)}(n)$ is won by player $2.$
Despite the interest of many mathematicians, little progress has been made on fully resolving ˜3 and conjectures 2 and 4, or the unbiased $p=q$ versions. At the end of our work, we highlight the underlying difficulties that may explain why this is so hard and remains unresolved.
In Section˜5, we consider the three games for Colex graphs. A Colex graph $\mathcal{C}(m)$, where $m=\binom{m_{1}}{2}+m_{2}$ with $0\leq m_{2}<m_{1}$, is the graph formed by adding $m_{2}$ edges between $K_{m_{1}}$ and an additional vertex (if $m_{2}>0$). See Fig.˜3 for examples where $(m_{1},m_{2})\in\{(4,3),(5,2)\}.$ As a class of graphs that smoothly extend cliques, this serves as a showcase to challenge certain intuitive expectations by exhibiting particular behaviours. The results were found by computer code (which implements the idea of Zermelo’s theorem [13]), explained in Section˜6.
2 Further results by strategy stealing
In this section, we present the proof of ˜5.
Proof of ˜5..
[…]
As a corollary of ˜5, together with the computer verification for $n\leq 8$, we know that ˜4 is true for $n\leq 9.$
When $4\mid n,$ the truth of ˜4 appears natural since we know that $s_{\operatorname{VC}}(G)\in\{0,2\}$, while the size and order are both even. Nevertheless, taking into account that $\operatorname{VC}(\mathcal{C}(12))\not=0$ (see Table˜1) while $\mathcal{C}(12)$ has even order and even size, a proof cannot be trivial. Also one might think about mirror strategies creating isomorphic coloured graphs on even size cliques $K_{n}.$ Nevertheless, already for $n\in\{4,5\}$ it can be checked that there is no winning strategy for the second player for the creation of isomorphic coloured graphs.
In [11], the authors proved that if Alice wins $\operatorname{Clique}(n)$ (something that conjecturally is only the case when $n=2$), then Bob wins $\operatorname{Clique}(n+3)$ for $n\geq R(5,5).$ This was later included on https://www.erdosproblems.com/778 (at that time) without the condition that $n$ has to be sufficiently large. Here we prove that it is indeed a valid statement for all $n$. This follows in essence from [11, Rem. 8] with the additional observation that $n\geq R(3,4)=9$ instead of $n\geq R(4,4)=18$ also works, and the computer check for $n\leq 8$ (checked with independent implementation for $n\leq 7$).
Proposition 9.
If Alice wins $\operatorname{Clique}(n)$ for $n\geq 8$, then Bob wins $\operatorname{Clique}(n+3).$ Bob wins $\operatorname{Clique}(n)$ for $3\leq n\leq 8.$
Proof.
Observation 10.
Proof.
Figure 1: The 55 possible situations after 33 edges have been coloured.
Bob colours a second edge, not incident with his first, to obtain one of the following three configurations (Fig.˜2). Note that his lower (the lower one drawn in Fig.˜2) blue edge is incident with both of the red edges. One can check each of those $3$ cases separately. Independently of what Alice does in her third turn, Bob can answer, ensuring a triangle $uv_{1}v_{2}$ with two out of three edges coloured blue (called $uv_{1}$ and $uv_{2}$), the subgraph spanned by $V\setminus\{u,v_{1},v_{2}\}$ containing exactly one coloured edge, which is blue. We show this by giving his action in three cases.
In the remaining of the game, Bob colours an edge spanned by $V\setminus\{u,v_{1},v_{2}\}$ whenever Alice coloured one there, using the strategy on the $K_{n}$ spanned by $V\setminus\{u,v_{1},v_{2}\}$ to win (as first player). If Alice coloured an other edge (incident with at least one of $\{u,v_{1},v_{2}\})$, Bob colours an uncoloured edge from a pair $\{v_{1}x,v_{2}x\}$ with $x\in V\setminus\{u,v_{1},v_{2}\}$ where the other edge is red, whenever possible, and otherwise a random edge. By strategy stealing, Bob can obtain a larger clique on the $K_{n}$ spanned by $V\setminus\{u,v_{1},v_{2}\}$ and thus not lose on $K_{n+3}$. For this, note that the largest clique on $K_{n+3}$ coloured by Alice is bounded by the largest one containing $v_{1}v_{2}$ (at most a $K_{4}$, since she could not colour three additional pairs in red) and the largest clique on $V\setminus\{u,v_{1},v_{2}\}$ plus one.
Figure 2: The 33 possible situations after Bob colours its second edge.
3 The biased games for large $n$
We answer the final question in the conclusion of [11], by noticing the relation with a result on the (degree) balancing game by Beck [2, Thm. 17.5].
Theorem 11.
When $p<q$ are integers, then $\Delta_{(p,q)}(n)$ and $\operatorname{VC}_{(p,q)}(n)$ are both player $2$ wins provided that $n$ is sufficiently large.
Proof.
Bob can play as the first player for the balancing game from [2, Thm. 17.5], neglecting the first $p$ edges from Alice. Consequently, he can ensure that every vertex has blue degree $\frac{q+o_{n}(1)}{p+q}n$ (recall that $p=o(n)$ as well). For $n$ sufficiently large, all of these degrees are larger than $\frac{n}{2}$ and so Bob wins both the maximum degree game and the vertex capturing game, since $\delta_{B}>\frac{n}{2}>\Delta_{A}$. ∎
It can be observed that for $(p,q)=(1,2),$ $n\geq 3$ is sufficient.
Proposition 12.
The games $\Delta_{(1,2)}(n)$ and $\operatorname{VC}_{(1,2)}(n)$ are both player $2$ wins provided that $n\geq 3.$
Proof.
Write $n=3k+r$, where $k\geq 1$ and $0\leq r\leq 2.$ Let the vertices be partitioned in $U=\{u_{1},u_{2},\ldots,u_{k}\},V=\{v_{1},v_{2},\ldots,v_{k}\}$ and $W=\{w_{1},\ldots,w_{k}\}$, and also $x$ if $r=1$ or $x$ and $y$ if $r=2$.
Bob colours the two other vertices for triples of the following form:
$\{u_{i}u_{j},v_{i}v_{j},w_{i}w_{j}\},\{u_{i}v_{i},u_{i}w_{i},v_{i}w_{i}\},\{u_{i}v_{j},v_{i}w_{j},w_{i}v_{j}\},\{xu_{i},xv_{i},xw_{i}\},\{yu_{i},yv_{i},yw_{i}\},$
where $1\leq i,j\leq k.$ In this case the number of blue incident edges to (at least one) of $u_{i},v_{i},w_{i}$ is double the number of red such edges. As such, $\Delta_{A}<\Delta_{B}$ and Bob captured at least $k$ vertices more than Alice did. ∎
Next, we consider regimes in which $\operatorname{Clique}_{(p,q)}(n)$ is won by the second player.
By applying [4, Thm. 2], we get a general improvement for the general biased clique game. We recall the latter statement for convenience, formulated slightly different (the way applied in [4, Thm. 1]).
Theorem 13 ([4]).
Let $G$ be a graph of order $n$ and size bounded by $\frac{\binom{n}{2}}{k}$. Then $\max\{\omega(G),\alpha(G)\}>c_{4}\frac{k}{\log k}\log n.$
Theorem 14.
There is a constant $c$ such that if $q>cp^{2}$, then for every $n$ sufficiently large, $\operatorname{Clique}_{(p,q)}(n)$ is won by player $2.$
Proof.
Let $c=\frac{2}{c_{4}},$ where $c_{4}$ is as in ˜13. Then for $n$ sufficiently large, the ratio of edges coloured by the first player is bounded by $p+\frac{p}{p+q}\binom{n}{2}<\frac{\binom{n}{2}}{k}$ for $q+1\geq k\geq\frac{q}{p}>\frac{2}{c_{4}}p.$ By ˜13, the graph of at least one of the two players contains a clique of order no less than $c_{4}\frac{k}{\log k}\log n.$
By [2, Thm. 20.1] on the other hand, the second player can force the largest clique induced by the colour of the first player to be bounded by $\frac{2p}{\log(q+1)}\log n.$
Since $c_{4}\frac{k}{\log k}>\frac{2p}{\log(q+1)},$ we conclude the second player wins. ∎
For further (conditional) improvement, we consider Maker-Breaker games, in which Maker tries to obtain a clique $K_{q}$ in its colour, and Breaker tries to prevent that.
Let $f_{N}(m,b)$ denote the largest $q$ such that Maker can occupy a $K_{q}$ in the $(m:b)$ game (Maker has $m$ edges and Breaker has $b$ edges) played on $K_{N}$. The value of $f_{N}(1,1)\sim 2\log_{2}N$ is known exactly by Beck [2, Thm. 6.4]. Beck [2, Thm. 30.2] also gave lower bounds for $f_{N}(m,1)$ of the form $2(1+o(1))\log_{(m+1)/m}(N)$ and Gebauer [5, Thm. 1.3] gave a lower bound on $f_{N}(m,b)$, $(1+o(1))m\log_{b+1}(N)$.
Motivated by these lower bounds, we state the following conjecture.
Conjecture 15.
For fixed $b,m$, for $N$ sufficiently large $f_{N}(m,b+1)<f_{N}(m,b)<f_{N}(m+1,b)$ and the differences tend to infinity as $N$ goes to infinity.
The first inequality for $m=1$ would follow from the unsolved final conjecture by Bednarska and Łuczak [3].
This strict monotonicity immediately implies that ˜14 can be further improved.
Theorem 16.
If ˜15 is true, then for fixed $p$ and $n$ sufficiently large, $\operatorname{Clique}_{(p,2p+1)}(n)$ is won by player $2.$
Proof.
The second player uses $p$ edges as Breaker for the formation of a clique by the first player, and $p+1$ edges to form a large clique itself. The first $p$ edges of player $1$ can be omitted in the building process of the second player, and if his own edges are already placed at optimal places, other edges can be put randomly. We deduce that $\omega_{A}\leq f_{N}(p,p)$ and $\omega_{B}\geq f_{N}(p+1,p)-p.$ By assumption, for $N$ sufficiently large this implies $\omega_{B}>\omega_{A}$. ∎
By the results from Beck, this in particular implies that the second player wins $\operatorname{Clique}_{(1,3)}(n)$ provided that $n$ is sufficiently large. In Section˜4, we prove this for all $n\geq 4$.
It is also natural to conjecture the following sharp version for the game.
Conjecture 17.
Provided that $n$ is sufficiently large, the $\operatorname{Clique}_{(p,p+1)}(n)$ is won by player $2.$
4 The biased clique game $\omega_{1,3}$
There are multiple elementary proofs possible when $n$ is even. We provide the one solution we could modify for the odd case.
Theorem 18.
For every $n\geq 4$ even, $\operatorname{Clique}_{(1,3)}(n)$ is won by player $2.$
Proof.
We will consider a turn as one move (colouring an edge) by Alice and (at most) $3$ moves of Bob. Since Bob can also play some of the remaining edges randomly, the win for Bob will be kept, so we present the strategy for Bob in this simpler way.
The $n=2k$ vertices are initially unlabeled vertices. During the game, a vertex is labeled as soon as at least one of its incident edges is coloured, or when there are only two vertices without coloured incident edges, in which case both are labeled.
Initial play Here we assume that at the beginning of the turn, at least $4$ vertices are unlabeled. The $2m$ labeled vertices will have indices which appear in pairs $(v_{i},v_{i+k})$, where $1\leq i\leq m$. The edges $v_{i}v_{i+k}$, will be coloured by Bob. If an edge $v_{i}v_{j}$ is coloured by Alice, the symmetric edge $v_{i+k}v_{j+k}$ will have been coloured by Bob. Here indices can be considered modulo $2k$ throughout.
Depending on the edge Alice chooses, Bob reacts as follows.
  • If Alice colours an edge between two unlabeled vertices, we will label them as $v_{i},v_{j}$ (where $i=m+1,j=m+2$). Bob picks two unlabeled vertices, calls them $v_{i+k},v_{j+k}$, and colours the edges $v_{i}v_{i+k},v_{i+k}v_{j+k},v_{j}v_{j+k}$.
  • If Alice colours an edge whose endvertices contain a labeled one, $v_{i}$, and one unlabeled one, which we call $v_{m+1}.$ Bob takes an other unlabeled vertex, $v_{m+1+k},$ and colours $v_{i+k}v_{m+1+k}$ and $v_{m+1}v_{m+1+k}$.
  • If Alice colours an edge $v_{i}v_{j}$ whose end vertices are already labeled, Bob colours $v_{i}v_{j+k},v_{i+k}v_{j+k}$ and $v_{j}v_{i+k}$ (some of these edges may have been coloured before by Bob if the vertices $v_{i},v_{j}$ were labeled in the same turn).
If two unlabeled vertices remain at the end of the turn, we label them by $x$ and $y$, and we go to the final play.
Finishing play
  • If Alice colours $xy$, Bob colours no edges.
  • If Alice colours $xv_{i}$, then Bob colours $xv_{i+k},yv_{i}$ and $yv_{i+k}$.
  • If Alice colours $yv_{i}$, then Bob colours $xv_{i},xv_{i+k}$ and $yv_{i+k}$.
  • If Alice colours an edge $v_{i}v_{j}$ whose end vertices are already labeled, then Bob responds as during the initial play; that is, he colours $v_{i}v_{j+k},v_{i+k}v_{j+k}$ and $v_{j}v_{i+k}$ (some of these edges may have been coloured before by Bob if the vertices $v_{i},v_{j}$ were labeled in the same turn).
Having sketched a winning strategy for Bob, we prove that Bob wins.
Consider a largest clique induced by all vertices different from $x$ and $y$ for which all edges are coloured by Alice. Up to shifting and relabeling opposite vertices, we denote it $v_{1}v_{2}\ldots v_{r}.$ This clique has order $r\geq 2.$ In that case, by the strategy of Bob, he coloured the clique $v_{k+1}v_{k+2}\ldots v_{k+r}$ containing all symmetric edges.
There are at most r1r-1 edges among $G[v_{1},v_{2},\ldots,v_{r}]$ which have been coloured by Alice at a moment at least one of its endvertices was unlabeled. For each of these edges $v_{i}v_{j}$, the clique $G[v_{i},v_{j},v_{i+k},v_{j+k}]\cong K_{4}$ contains two edges coloured by Alice. If $v_{i}v_{j}$ was already labeled before Alice coloured that edge, it is the only edge within $G[v_{i},v_{j},v_{i+k},v_{j+k}]$ which is coloured by her. This implies that the bipartite graph spanned by the partition classes $V_{1}=\{v_{1},v_{2},\ldots,v_{r}\}$ and $V_{2}=\{v_{k+1},v_{k+2},\ldots,v_{k+r}\}$ has at most r1r-1 blue edges. This implies that there is a vertex $v_{i}$ in $V_{1}$ such that $G[v_{i},V_{2}]$ is fully coloured by Bob. That is, $G[V_{2}\cup\{v_{i}\}]$ is a clique of order $r+1$ fully coloured by Bob.
By Bob’s strategy, Alice has no triangle containing the edge $xy$ that is fully coloured. If Alice has a large fully coloured clique containing $x$, then it must be of the form $V_{1}\cup\{x\}$; analogously, if it contains $y$, it must be of the form $V_{1}\cup\{y\}$. But then the clique spanned by $V_{2}\cup\{v_{i},y\}$ (resp. $V_{2}\cup\{v_{i},x\}$) is fully coloured by Bob. ∎
Theorem 19.
For every $n\geq 5$ be odd, $\operatorname{Clique}_{(1,3)}(n)$ is won by player $2.$
Proof.
The setup and initial play proceed as in ˜18, up to the point where either three unlabeled vertices remain and Alice colours an edge incident to at least one of them, or five unlabeled vertices remain and Alice begins colouring an edge between two of these. The labeled vertices (possibly after relabeling), can be $L=\{v_{i}\mid 1\leq i\leq 2k\}$ where $v_{i}$ and $v_{i+k}$ form a pair of opposite vertices. Here $n=2k+3$ or $n=2k+5$.
When $n=2k+3$, we call the the three remaining unlabeled vertices $a,b,c.$
When $n=2k+5$, we name the end vertices of the edge coloured by Alice $x,y$. We label the other three by $a,b,c.$
We present the final play in each of the following three situations.
Case 1: $n=2k+3$ and Alice colours $av_{\ell}$
Bob colours the edges of abcabc.
From now on;
  • If Alice colours an edge in $\{a,b,c\}\times\{v_{\ell},v_{\ell+k}\}$, then Bob can colour edges among $\{a,b,c\}\times\{v_{\ell},v_{\ell+k}\}$ (but does not even have to).
  • If Alice colours an edge $v_{i}v_{j}\in G[L]$, Bob colours $v_{i}v_{j+k},v_{i+k}v_{j+k}$ and $v_{j}v_{i+k}$ (some of these edges may have been coloured before by Bob if the vertices $v_{i},v_{j}$ were labeled in the same turn).
  • If Alice colours an edge in $dv_{i}$ for some $d\in\{a,b,c\}$, then Bob colours $av_{i+k},bv_{i+k}$ and $cv_{i+k}$. If those three were already coloured, Bob colours the uncoloured edge among $av_{i},bv_{i}$ and $cv_{i}$.
If a largest clique of Alice is contained in $G[L],$ Bob has a larger clique within $G[L],$ as shown before in the proof of ˜18.
If the largest clique contains one vertex from $\{a,b,c\}$ and w.l.o.g. $V_{1}=\{v_{1},v_{2},\ldots,v_{r}\}$ ($r$ vertices from $L$.
Let $V_{2}=\{v_{k+1},v_{k+2},\ldots,v_{k+r}\}$.
Bob has a clique spanned by the vertices $V_{2}\setminus\{v_{\ell},v_{\ell+k}\}\cup\{a,b,c\}$, so its clique number is at least $r+2.$
Case 2: $n=2k+3$ and Alice colours abab
Bob reacts by colouring the other two edges of abcabc.
From now on;
  • If Alice colours an edge $av_{i}$, then Bob colours $bv_{i},bv_{i+k}$ and $cv_{i+k}.$ Symmetrically, if Alice colours an edge $bv_{i}$, then Bob colours $av_{i},av_{i+k}$ and $cv_{i+k}.$
  • If Alice colours an edge in $cv_{i}$, then Bob colours $av_{i+k},bv_{i+k}$ and $cv_{i+k}.$
  • If Alice colours an edge $v_{i}v_{j}\in G[L]$, Bob colours $v_{i}v_{j+k},v_{i+k}v_{j+k}$ and $v_{j}v_{i+k}$ (some of these edges may have been coloured before by Bob if the vertices $v_{i},v_{j}$ were labeled in the same turn).
If a largest clique of Alice is contained in $G[L],$ Bob wins as before.
If a largest clique of Alice is of the form $\{a,v_{1},v_{2},\ldots,v_{r}\}$, then Bob has a clique spanned by $\{b,c,v_{k+1},v_{k+2},\ldots,v_{k+r}\}$. Similarly, $\{b,v_{1},v_{2},\ldots,v_{r}\}$ or $\{c,v_{1},v_{2},\ldots,v_{r}\}$ would be a smaller clique than $\{a,c,v_{k+1},v_{k+2},\ldots,v_{k+r}\}$.
Case 3: $n=2k+5$ and Alice colours xyxy
Bob reacts by colouring the edges of abcabc.
From now on;
  • If Alice colours an edge in $av_{i},bv_{i}$ or $cv_{i}$, then Bob colours $av_{i+k},bv_{i+k}$ and $cv_{i+k}.$
  • If Alice colours $xv_{i}$ or $yv_{i},$ Bob colours the three remaining edges in $\{x,y\}\times\{v_{i},v_{i+k}\}.$
  • If Alice colours w.l.o.g. $ax$, Bob colours $ay,by$ and $cy.$ (the $6$ edges are all equivalent) If later, Alice colours $bx$ or $cx$, Bob colours the other edge among those two.
  • If Alice colours an edge $v_{i}v_{j}\in G[L]$, Bob colours $v_{i}v_{j+k},v_{i+k}v_{j+k}$ and $v_{j}v_{i+k}$ (some of these edges may have been coloured before by Bob if the vertices $v_{i},v_{j}$ were labeled in the same turn).
If a maximum clique from Alice contains at most one vertex from $\{a,b,c,x,y\},$ then Bob has a larger clique as before.
The only clique of Alice containing $xy$ is $xy$ itself.
If $\{a,x,v_{1},v_{2},\ldots,v_{r}\}$ is a maximum clique from Alice, then $\{b,c,y,v_{k+1},v_{k+2},\ldots,v_{k+r}\}$ is a monochromatic clique coloured by Bob. In all cases, the maximum clique coloured by Bob is strictly larger than the maximum clique coloured by Alice. ∎